# LCEN: A Novel Feature Selection Algorithm for Nonlinear, Interpretable Machine Learning Models

### v1

@@ Abstract
Interpretable architectures can have advantages over black-box architectures, and interpretability is essential for the application of machine learning in critical settings, such as aviation or medicine. However, the simplest, most commonly used interpretable architectures (such as LASSO or EN) are limited to linear predictions and have poor feature selection capabilities. In this work, we introduce the LASSO-Clip-EN (LCEN) algorithm for the creation of nonlinear, interpretable machine learning models. LCEN is tested on a wide variety of artificial and empirical datasets, creating more accurate, sparser models than other commonly used architectures. These experiments reveal that LCEN is robust against many issues typically present in datasets and modeling, including noise, multicollinearity, data scarcity, and hyperparameter variance. LCEN is also able to rediscover multiple physical laws from empirical data and, for processes with no known physical laws, LCEN achieves better results than many other dense and sparse methods – including using 10.8 times fewer features than dense methods and 8.1 times fewer features than EN on one dataset, and is comparable to an ANN on another dataset.
Machine Learning, Variable Selection, Elastic Net, Interpretable Machine Learning
@@ 1 Introduction
Statistical models are powerful tools to explain, predict, or describe natural phenomena (Shmueli, 2010). They connect independent variables (also called “inputs” or “features”) to dependent variables (also called “outputs” or “labels”) to test causal hypotheses, predict novel outputs from known inputs, or summarizing the data structure (Shmueli, 2010). Many model architectures exist, including linear, ensemble-based, and deep learning models. Complex architectures are claimed to have greater capability to model phenomena due to their lower bias, but their intricate and numerous mathematical transformations prevent humans from understanding how an output was predicted by a model, or the relative or absolute importance of the inputs. Moreover, a lack of transparency may prevent the model from being trusted in critical or sensitive applications (Hong et al., 2020).
@@
As summarized by Ribeiro et al. (2016), there are two main methods to increase interpretability: the use of model-agnostic algorithms, which extract interpretable explanations a posteriori and work for any architecture, or the direct use of interpretable architectures. Interpretable architectures include “decision trees, rules, additive models, attention-based networks, and sparse linear models” (Ribeiro et al., 2016). As elaborated in the review of Rudin (2019), interpretable architectures can have many advantages over black-box or a posteriori explanations, including the ability to assist researchers in refining the model and data, or better highlighting scenarios in which the model fails or lacks robustness. Special attention should be given to sparse models, which identify the most important features, can make the model more robust to variations in the input data, and significantly improve the model’s interpretability (Rudin, 2019). At the same time, even a linear model or decision tree/rules can become unwieldy and uninterpretable if hundreds or thousands of coefficients or rules are present.
@@
Feature selection is the process of selecting the most important features in a model to increase its robustness or interpretability. Many criteria for feature selection exist (Heinze et al., 2018), including significance based on p-values (using a univariate, iterative/stepwise, or global method), using information criteria (such as the AIC (Akaike, 1974) and BIC (Schwarz, 1978)), using penalties (such as in LASSO (Santosa & Symes, 1986) and elastic net [EN] (Zou & Hastie, 2005)), criteria based on changes in estimates, and expert knowledge. While no method is superior for all problems, different works have evaluated and criticized these criteria. For example, stepwise regression is one of the most commonly used methods in many fields thanks to its computational simplicity and ease of understanding (Whittingham et al., 2006; Heinze et al., 2018; Smith, 2018). However, stepwise regression is prone to ignoring features with causal effects, including irrelevant features, generating excessively small confidence intervals, and producing incorrect/biased parameters (Whittingham et al., 2006; Smith, 2018). LASSO is simple and computationally cheap, and has performed well for some problems (Hebiri & Lederer, 2013; Tian et al., 2015; Pavlou et al., 2016), but can overselect irrelevant variables, can select only as many features as there are samples, and does not handle multicollinear data well [^1] This last point is somewhat controversial in the literature, see Hebiri & Lederer (2013) and Dalalyan et al. (2017), for example. (Heinze et al., 2018; Zou & Hastie, 2005).
@@
Most feature selection methods apply only in linear contexts (or have been applied primarily in linear contexts). For example, the only sparse models referenced in the highly cited review by Ribeiro et al. (2016) are linear. The main sparse architectures (LASSO, EN, and their variants) are linear regressors. Later works consider sparse nonlinear models, to address the fact that many natural and industrial processes are nonlinear. McConaghy (2011), Brewick et al. (2017), and Sun & Braatz (2020) defined sets of features consisting of polynomials (all works), interactions (all works), and/or non-polynomials (McConaghy, 2011; Sun & Braatz, 2020). ALVEN, the model architecture from Sun & Braatz (2020), uses an f-test for each feature (including the expanded set of features) to determine whether to keep a feature in the final EN model. However, this f-test has very poor feature selectivity, as nearly all features are selected when traditional values of $\alpha$ ($0.001\leq\alpha\leq 0.05$) are used. Furthermore, the ordering of the features with respect to their p-values does not follow their relevance, as many irrelevant features are among those with the lowest p-values, and relevant features can be among those with the highest p-values (including p $\gg 0.05$).
@@
To create nonlinear, interpretable machine learning models with high predictive and descriptive power, we propose the LASSO-Clip-EN (LCEN) algorithm. This algorithm generates an expanded set of features (such as in ALVEN) and performs feature selection and model fitting. The algorithm is tested on artificial and empirical data, successfully rediscovering physical laws from data belonging to multiple different areas of knowledge with errors < 2% on the coefficients, a value within the empirical noise of the datasets. On datasets from processes whose underlying physical laws are not yet known, LCEN attains lower RMSEs than sparse and dense methods and simultaneously leads to sparser models than alternative architectures.
@@ 2.1 Datasets
Multiple datasets are used to test the performance of the LCEN algorithm (summarized in Table A1). These datasets can be divided into three categories: artificial data [“Linear, 5-variable polynomial”, “Multicollinear data”, “Relativistic energy”, and “4th-degree, univariate polynomial”], empirical data from processes with known physical laws [“CARMENES star data”, “Kepler’s 3rd Law”, and “Newton’s Law of Cooling”], and empirical data from processes with no known physical laws [“Diesel Freezing Point”, “Abalone”, and “Concrete Compressive Strength”]. The artificial data generated by us are used for an initial assessment of the LCEN algorithm and to investigate how properties of the data, such as noise or data range, affect its feature selection capabilities. Empirical data from processes with known physical laws are used to verify whether the LCEN algorithm can rediscover known physical laws using data with real properties. Empirical data from processes with no known physical laws are used to compare the performance of the LCEN algorithm against other linear and nonlinear models, including deep learning models. Further description of these datasets and how the artificial datasets are generated is available in Section A1.
All models tested in this work, with the exception of the model for the “Newton’s Law of Cooling” dataset, had their hyperparameters selected by 5-fold cross-validation. The separation between training and testing sets varied depending on the dataset. None of the artificial datasets or datasets containing empirical data from processes with known physical laws have a separate test set, as they are used to investigate the capability of the LCEN algorithm to select correct features (which occurs based on the training set). For the “Diesel freezing point” dataset, 30% of the dataset was randomly separated to form the test set. For the “Abalone” dataset, the last 1,044 entries (25%) were used as the test set as per Waugh (1995) and Clark et al. (1996). For the “Concrete Compressive Strength” dataset, 25% of the dataset was randomly separated to form the test set as per Yeh (1998).
@@ 2.2 Algorithm
The LCEN algorithm has five hyperparameters: alpha, which determines the regularization strength (as in the LASSO, EN, and similar algorithms); l1_ratio, which determines how much of the regularization of the EN step depends on the 1-norm as opposed to the 2-norm (as in the EN algorithm); degree, which determines the maximum degree for the basis expansion of the data (Table A2); lag, which determines the maximum number of previous time steps from which features are included (relevant only for dynamic models); and cutoff, which determines the minimum value a scaled parameter needs to have to not be eliminated during the clip step.
The LCEN algorithm (Algorithm 1) begins with the LASSO step, setting the l1_ratio to 1. Cross-validation on the training set is employed among all combinations of alpha, degree, and lag values. The values of degree and lag corresponding to the model with the lowest validation MSE are recorded, and parameters are obtained for this LASSO model. The next step in the LCEN algorithm is the clip step, in which parameters smaller than the cutoff are set to 0. Finally, the EN step involves cross-validation on the training set among all combinations of alpha and l1_ratio, using the values of degree and lag obtained in the LASSO step. A final, nonlinear, and interpretable model, whose coefficients are subjected to the clip step again, is returned.
Algorithm 1 LASSO-Clip-EN (LCEN)
  Input: data $X$ and $y$; lists of hyperparameters alpha, l1_ratio, degree, lag; hyperparameter cutoff # LASSO step
  Temporarily set l1_ratio$=1$.
  for each combination in (alpha, degree, lagdo
     Perform cross-validation.
  end for
  Obtain the best hyperparameters from the above cross-validation.
  Obtain parameters with these hyperparameters.
  Record the best degree and lag hyperparameters. # Clip step
  Remove all parameters < cutoff. # EN step
  Restore l1_ratio to its original list of values.
  for each combination in (alpha, l1_ratiodo
     Perform cross-validation.
  end for
  Obtain the best hyperparameters from the above cross-validation.
  Obtain parameters with these hyperparameters. # Clip step II
  Remove all parameters < cutoff.
  return
@@ 3.1 Artificial data highlight LCEN’s robustness to noise, multicollinearity, and hyperparameter variance
The first dataset used to validate the LCEN algorithm is a simple linear polynomial with five independent variables (“Linear, 5-variable polynomial”). This polynomial is of the form $y=-2.8X_{0}-2.7X_{1}-5.3X_{2}+4.3X_{3}+9.0X_{4}+\epsilon$. Datasets were created with increasing noise levels to determine how LCEN performs. For noise levels up to 119%, LCEN returned the correct polynomial among nonlinear and non-polynomial features with degree $\leq 3$ (Table 1). A noise level of 238%, the next level tested, led to the selection of one additional variable (${1}/{X_{4}}$), but the linear features had coefficients close to the ground truth. Furthermore, some of the coefficients did not display any appreciable change despite the increasing noise. This analysis highlights how LCEN is robust to noise in a scenario where all features are independent, even if the hyperparameters allow for extraneous features to potentially be selected by the algorithm.
@@
Table 1: Coefficient values and MSEs to the ground truth for the “Linear, 5-variable polynomial” dataset at different noise levels. The coefficients are ordered such that the $n$th value corresponds to the coefficient of $X_{n}$.
Noise Coefficients MSE 0% -2.80, -2.70, -5.30, 4.30, 9.00 9.6×$\times 10^{-14}$ 59.6% -2.82, -2.70, -5.27, 4.30, 9.01 2.2×$\times 10^{-4}$ 119% -2.83, -2.70, -5.24, 4.30, 9.01 8.7×$\times 10^{-4}$ 238% -2.85, -2.71, -5.20, 4.30, 8.73 1.7×$\times 10^{-2}$
@@
The first step of the LCEN algorithm uses LASSO, which has been claimed to underperform with multicollinear data (Heinze et al., 2018; Zou & Hastie, 2005). Therefore, tests using multicollinear data are done next. The goal is to verify whether LCEN can successfully determine the presence of two different but correlated variables, as LASSO prefers to select only one variable in this scenario (Zou & Hastie, 2005). Noise $\epsilon_{1}$, at different levels, was added to the $X_{0}$ variable to create a correlated variable $X_{1}$. A second noise $\epsilon_{2}$, also at different levels, was added to the final $y$ data. When $\epsilon_{1}=0$, both variables are equal and separation is not possible. However, at other $\epsilon_{1}$ values, the LCEN algorithm is very successful at identifying that two relevant variables exist and assigning correct coefficients to them (Figure 1). Specifically, when the noise level $\epsilon_{1}$ associated with the $X$ data (which indicates how different the $X$ variables are, as highlighted by the variance inflation factors [VIFs] in Fig. 1) is greater than the noise level $\epsilon_{2}$ associated with the $y$ data, LCEN can separate both variables with coefficient errors $\leq 5\%$. When both noise levels are similar, LCEN can separate both variables with coefficient errors between 5% and 10%. The $X$ data used in this experiment has very high multicollinearity (as shown by the VIFs); real data will typically have lower VIFs and thus be easier to separate using LCEN.
Figure 1: Output of the LCEN model at different $X$-data noise levels $\epsilon_{1}$ and $y$-data noise levels $\epsilon_{2}$. Bright red squares indicate both variables were selected and their coefficients had errors $\leq 5\%$. Light red squares indicate that both variables were selected and their coefficients had $5\%<\text{errors}\leq 10\%$. Squares in white indicate that both variables were selected and their coefficients had $10\%<\text{errors}\leq 20\%$. Light blue squares indicate that both variables were selected and their coefficients had errors > 20%. Bright blue squares indicate that only one of the variables was selected.
@@
Next, a more complex equation is used to further validate LCEN. The “Relativistic energy” data contain mass and velocity values used to calculate $E^{2}=c^{4}m^{2}+c^{2}m^{2}v^{2}$. As before, datasets with increasing noise levels are created. The degree hyperparameter is allowed to vary between 1 and 6 in this experiment. LCEN selected only relevant features for a noise level $\leq 15\%$, and the coefficients were equal to the ground truth (Table 2). For a noise level of 20%, LCEN selected the two correct features, yet it also selected nonzero coefficients for $m$ and $m^{3}$. This error led to our hypothesizing that it is challenging to distinguish among $m$, $m^{2}$, and $m^{3}$ due to the low range of the data. Thus, another dataset with the same properties but a larger range of values for $m$ is created. LCEN performed better on this dataset, selecting only relevant features for all noise levels $\leq 30\%$ (the highest value tested) and having much lower errors in the estimated coefficients (Table 2). These experiments further highlight the robustness of LCEN and show how the range of the data can affect predictions.
[…]
Noise Coefficients RMSE (%) 0% 8.077×$\times 10^{33}$, 8.987×$\times 10^{16}$ 0.013, 0.009 5% 8.081×$\times 10^{33}$, 8.969×$\times 10^{16}$ 0.043, 0.206 10% 8.085×$\times 10^{33}$, 8.951×$\times 10^{16}$ 0.097, 0.410 15% 8.089×$\times 10^{33}$, 8.935×$\times 10^{16}$ 0.139, 0.580 20% 5.579×$\times 10^{33}$, 8.921×$\times 10^{16}$ 30.94, 0.740
Noise Coefficients RMSE (%) 0% 8.078×$\times 10^{33}$, 8.988×$\times 10^{16}$ 0.002, 0.008 5% 8.077×$\times 10^{33}$, 8.987×$\times 10^{16}$ 0.014, 0.011 10% 8.077×$\times 10^{33}$, 8.985×$\times 10^{16}$ 0.055, 0.024 15% 8.068×$\times 10^{33}$, 8.983×$\times 10^{16}$ 0.116, 0.050 20% 8.070×$\times 10^{33}$, 8.981×$\times 10^{16}$ 0.093, 0.068 30% 8.061×$\times 10^{33}$, 8.979×$\times 10^{16}$ 0.210, 0.099
@@
Finally, LCEN is compared with the feature selection algorithm in ALVEN (Sun & Braatz, 2020), which uses the same basis function expansion, but uses f-tests for feature selection. The “4th-degree, univariate polynomial” dataset is created as per Sun & Braatz (2021), such that $y=X+0.5X^{2}+0.1X^{3}+0.05X^{4}+\epsilon$, 30 $X$ points are available for training, and 1,000 $X$ points are available for testing. These conditions simulate the scarcity of data potentially present in real datasets while ensuring test errors can be predicted with high confidence. Sun & Braatz (2021) created four types of ALVEN models for this prediction: one that always uses degree = 4 (“unbiased model”), one that always uses degree = 2 (“biased model”), one that selects a degree between 1 and 10 based on cross-validation (“cv”), and one that selects a degree equal to 2 or 4 based on cross-validation (“cv limited order”). Sun & Braatz (2021) noted that the degree 4 “unbiased model” was the best at low noise levels, but its error quickly increases, leading to the degree 2 “biased model” becoming the best for noise levels > 75 (Figure 3 of Sun & Braatz (2021); reproduced with permission here as the left subfigure of Figure 2). The model with degree equal to 2 or 4 “cv limited order” was typically very close in performance to the best model at all noise levels, whereas the model with a degree between 1 and 10 “cv” had lower performance. Sun & Braatz (2021) explains these observations by the bias-variance tradeoff: at low noise levels, it is preferable to have the model follow the ground truth as closely as possible; thus, the degree 4 “unbiased model” was the best. However, once the noise level is sufficiently high, it becomes impossible to obtain enough signal within the data to compensate for the additional degrees of freedom (variance) in a 4th degree model; thus, the degree 2 “biased model” becomes the best. The degree between 1 and 10 “cv” model had lower performance due to its greater hyperparameter variance, and the degree equal to 2 or 4 “cv limited order” model struck a balance between the “unbiased model” and the “biased model”.

### v2

@@ Abstract
Interpretable architectures can have advantages over black-box architectures, and interpretability is essential for the application of machine learning in critical settings, such as aviation or medicine. However, the simplest, most commonly used interpretable architectures, such as LASSO or elastic net (EN), are limited to linear predictions and have poor feature selection capabilities. In this work, we introduce the LASSO-Clip-EN (LCEN) algorithm for the creation of nonlinear, interpretable machine learning models. LCEN is tested on a wide variety of artificial and empirical datasets, frequently creating more accurate, sparser models than other architectures, including those for building sparse, nonlinear models. LCEN is robust against many issues typically present in datasets and modeling, including noise, multicollinearity, data scarcity, and hyperparameter variance. LCEN is also able to rediscover multiple physical laws from empirical data and, for processes with no known physical laws, LCEN achieves better results than many other dense and sparse methods – including using 10.8-fold fewer features than dense methods and 8.1-fold fewer features than EN on one dataset, and is comparable to or better than ANNs on multiple datasets.
@@ 1 Introduction
Statistical models are powerful tools to explain, predict, or describe natural phenomena [1]. They connect independent variables (also called “inputs” or “features”) to dependent variables (also called “outputs” or “labels”) to test causal hypotheses, predict novel outputs from known inputs, or summarizing the data structure [1]. Many model architectures exist, including linear, ensemble-based, and deep learning models. Complex architectures are claimed to have greater capability to model phenomena due to their lower bias, but their intricate and numerous mathematical transformations prevent humans from understanding how an output was predicted by a model, or the relative or absolute importance of the inputs. Moreover, a lack of transparency may prevent the model from being trusted in critical or sensitive applications [2].
@@
As summarized by Ref. [3], there are two main methods to increase interpretability: the use of model-agnostic algorithms, which extract interpretable explanations a posteriori and work for any architecture, or the direct use of interpretable architectures. Interpretable architectures include “decision trees, rules, additive models, attention-based networks, and sparse linear models” [3]. It should be noted that nonlinear models may also be made sparse, and even interpretable, as described later in this section and the rest of this work. As elaborated in the review of Ref. [4], interpretable architectures can have many advantages over black-box or a posteriori explanations, including the ability to assist researchers in refining the model and data, or better highlighting scenarios in which the model fails or lacks robustness. Special attention should be given to sparse models, which identify the most important features, can make the model more robust to variations in the input data, and can significantly improve the model’s interpretability if an interpretable architecture is used [4]. At the same time, even a linear model or decision tree/rules can become unwieldy and uninterpretable if hundreds or thousands of coefficients or rules are present.
@@
Feature selection is the process of selecting the most important features in a model to increase its robustness or interpretability. Many criteria for feature selection exist [5], including significance based on p-values (using a univariate, iterative/stepwise, or global method), using information criteria (such as the AIC [6] and BIC [7]), using penalties (such as in LASSO [8, 9] and elastic net [EN] [10]), criteria based on changes in estimates, and expert knowledge. More broadly, these methods can be classified as filter, wrapper, or integrated methods. While no method is superior for all problems, different works have evaluated and criticized these criteria. For example, stepwise regression is one of the most commonly used methods in many fields thanks to its computational simplicity and ease of understanding [11, 5, 12]. However, stepwise regression is prone to ignoring features with causal effects, including irrelevant features, generating excessively small confidence intervals, and producing incorrect/biased parameters [11, 12]. LASSO is simple and computationally cheap, and has performed well for some problems [13, 14, 15], but can overselect irrelevant variables, can select only as many features as there are samples, and does not handle multicollinear data well [5, 10]. [^1] This last point is somewhat controversial in the literature, see Refs. [13, 16], for example.
@@
Originally, most feature selection methods applied only in linear contexts (or have been applied primarily in linear contexts). For example, the only sparse models referenced in the highly cited review by Ref. [3] are linear. The main sparse architectures (LASSO, EN, and their variants) are linear regressors. [^2] Ref. [17] has claimed that, if thousands of samples are available, LASSO can consistently select features even in nonlinear settings, although the coefficients may be distorted. However, this consistency is disputed by Ref. [18]. Later works consider sparse nonlinear models, to address the fact that many natural and industrial processes are nonlinear. Refs. [19, 20, 21] defined sets of features consisting of polynomials (all works), interactions (all works), and/or non-polynomials [19, 21]. ALVEN, the model architecture from Ref. [21], uses an F-test for each feature (including the expanded set of features) to determine whether to keep a feature in the final EN model, a filter approach. However, this F-test has very poor feature selectivity, as nearly all features are selected when traditional values of $\alpha$ ($0.001\leq\alpha\leq 0.05$) are used. Furthermore, the ordering of the features with respect to their p-values does not follow their relevance, as many irrelevant features are among those with the lowest p-values, and relevant features can be among those with the highest p-values (including p $\gg 0.05$). The work of Ref. [22] uses two forms of modified LASSO algorithms for nonlinear feature selection, achieving high sparsity with the datasets tested.
More recently, the L1 regularization has been applied to neural networks for nonlinear feature selection. In its simplest form, group LASSO is applied to zero all the outputs of some neurons (sparsifying the network and eliminating features when zeroing input-layer neurons) [18, 23, 24]. A slightly modified version of this algorithm applies the L1 penalty only to the input layer and includes a skip-connection between that layer and the output layer [25]. More complex applications of this method include the multi-modal neural networks of Ref. [26], the concrete autoencoders of Ref. [27], and the teacher-student network of Ref. [28]. The first two are notable for being at least partially unsupervised methods, suggesting they can select the most relevant features for a given dataset no matter the task. Some works have also used approaches other than the L1 norm, such as the L0 norm [29]. While these models are powerful tools for feature selection, two considerable limitations are present: first, they do not provide any information on how the selected features are contributing to the final prediction, significantly limiting interpretability. A posteriori methods to extract this information have been found unreliable [4], even if useful. Second, these complex model architectures may take “shortcuts” to make apparently accurate predictions [30, 31]. However, these “shortcuts” are not really relevant to the task, preventing proper generalization and human interpretation.
@@
To create nonlinear, interpretable machine learning models with high predictive and descriptive power, we propose the LASSO-Clip-EN (LCEN) algorithm. This algorithm generates an expanded set of nonlinear features (such as in ALVEN) and performs feature selection and model fitting. This feature set expansion, together with the Clip step, provide LCEN with the ability to do nonlinear predictions. The algorithm is tested on artificial and empirical data, successfully rediscovering physical laws from data belonging to multiple different areas of knowledge with errors < 2% on the coefficients, a value within the empirical noise of the datasets. On datasets from processes whose underlying physical laws are not yet known, LCEN attains lower RMSEs than many sparse and dense methods, leads to sparser models than any other architecture tested, and simultaneously runs faster than most alternative architectures.
@@ 2.1 Datasets
Multiple datasets are used to test the performance of the LCEN algorithm (summarized in Table A1). These datasets can be divided into three categories: artificial data [“Linear, 5-variable polynomial”, “Multicollinear data”, “Relativistic energy”, and “4th-degree, univariate polynomial”], empirical data from processes with known physical laws [“CARMENES star data” and “Kepler’s 3rd Law”], and empirical data from processes with no known physical laws [“Diesel Freezing Point”, “Abalone”, “Concrete Compressive Strength”, “Boston housing”, and “GEFCom 2014”]. The artificial data generated by us are used for an initial assessment of the LCEN algorithm and to investigate how properties of the data, such as noise or data range, affect its feature selection capabilities. Empirical data from processes with known physical laws are used to verify whether the LCEN algorithm can rediscover known physical laws using data with real properties. Empirical data from processes with no known physical laws are used to compare the performance of the LCEN algorithm against other linear and nonlinear models, including deep learning models. Further description of these datasets and how the artificial datasets are generated is available in Section A2.
All models tested in this work had their hyperparameters selected by 5-fold cross-validation. The separation between training and testing sets varied depending on the dataset. None of the artificial datasets or datasets containing empirical data from processes with known physical laws have a separate test set, as they are used to investigate the capability of the LCEN algorithm to select correct features (which occurs based on the training set). For the “Diesel freezing point” dataset, 30% of the dataset was randomly separated to form the test set. For the “Abalone” dataset, the last 1,044 entries (25%) were used as the test set as per Refs. [32, 33]. For the “Concrete Compressive Strength” dataset, 25% of the dataset was randomly separated to form the test set as per Ref. [34]. For the “Boston housing” dataset, 20% of the dataset was randomly separated to form the test set. For the “GEFCom 2014” dataset, the data from task 1 were used as the training set and all data from tasks 2–15 were used as the test set.
@@ 2.2 Algorithm
The LCEN algorithm has five hyperparameters: alpha, which determines the regularization strength (as in the LASSO, EN, and similar algorithms); l1_ratio, which determines how much of the regularization of the EN step depends on the 1-norm as opposed to the 2-norm (as in the EN algorithm); degree, which determines the maximum degree for the basis expansion of the data (Table A2); lag, which determines the maximum number of previous time steps from which features are included (relevant only for dynamic models); and cutoff, which determines the minimum value a scaled parameter needs to have to not be eliminated during the clip step. Details on the cross-validated hyperparameter values for all models are in Section A3.
The LCEN algorithm (formally typeset as Algorithm 1) begins with the LASSO step, setting the l1_ratio to 1. Cross-validation on the training set is employed among all combinations of alpha, degree, and lag values. The values of degree and lag corresponding to the model with the lowest validation MSE are recorded, and parameters are obtained for this LASSO model. The next step in the LCEN algorithm is the clip step, in which parameters smaller than the cutoff are set to 0. Finally, the EN step involves cross-validation on the training set among all combinations of alpha and l1_ratio, using the values of degree and lag obtained in the LASSO step. A final, nonlinear, and interpretable model, whose coefficients are subjected to the clip step again, is returned.
@@ 3.1 Artificial data highlight LCEN’s robustness to noise, multicollinearity, and hyperparameter variance
The first dataset used to validate the LCEN algorithm is a simple linear polynomial with five independent variables (“Linear, 5-variable polynomial”). This polynomial is of the form $y=-2.8X_{0}-2.7X_{1}-5.3X_{2}+4.3X_{3}+9.0X_{4}+\epsilon$. Datasets were created with increasing noise levels to determine how LCEN performs. For all noise levels, LCEN returned the correct polynomial among nonlinear and non-polynomial features with degree $\leq 3$ (Table 1). Furthermore, the coefficients did not display any appreciable change despite the increasing noise. This analysis highlights how LCEN is robust to noise in a scenario where all features are independent, even if the hyperparameters allow for extraneous features to potentially be selected by the algorithm.
@@
Table 1: Predicted coefficients and RMSEs to the true coefficients for the “Linear, 5-variable polynomial” dataset at different noise levels. The coefficients are ordered such that the $n$th value corresponds to the coefficient of $X_{n}$.
Noise Coefficients RMSE 0% -2.80, -2.70, -5.30, 4.30, 9.00 3.1×$\times 10^{-7}$ 59.6% -2.82, -2.70, -5.27, 4.30, 9.01 1.5×$\times 10^{-2}$ 119% -2.83, -2.70, -5.24, 4.30, 9.01 3.0×$\times 10^{-2}$ 238% -2.86, -2.70, -5.19, 4.29, 9.03 5.9×$\times 10^{-2}$
@@
The first step of the LCEN algorithm uses LASSO, which has been claimed to underperform with multicollinear data [5, 10]. Therefore, tests using multicollinear data are done next. The goal is to verify whether LCEN can successfully determine the presence of two different but correlated variables, as LASSO prefers to select only one variable in this scenario [10]. Noise $\epsilon_{1}$, at different levels, was added to the $X_{0}$ variable to create a correlated variable $X_{1}$. A second noise $\epsilon_{2}$, also at different levels, was added to the final $y$ data. When $\epsilon_{1}=0$, both variables are equal and separation is not possible. However, at other $\epsilon_{1}$ values, the LCEN algorithm is very successful at identifying that two relevant variables exist and assigning correct coefficients to them (Figure A1). Specifically, when the noise level $\epsilon_{1}$ associated with the $X$ data (which indicates how different the $X$ variables are, as highlighted by the variance inflation factors [VIFs] in Fig. A1) is greater than the noise level $\epsilon_{2}$ associated with the $y$ data, LCEN can separate both variables with coefficient errors $\leq 5\%$. When both noise levels are similar, LCEN can separate both variables with coefficient errors between 5% and 10%. The $X$ data used in this experiment has very high multicollinearity (as shown by the VIFs); real data will typically have lower VIFs and thus be easier to separate using LCEN.
@@
Next, a more complex equation is used to further validate LCEN. The “Relativistic energy” data contain mass and velocity values used to calculate $E^{2}=c^{4}m^{2}+c^{2}m^{2}v^{2}$. As before, datasets with increasing noise levels are created. The degree hyperparameter is allowed to vary between 1 and 6 in this experiment. LCEN selected only relevant features for all noise levels tested ($\leq 20\%$), and the coefficients were typically equal to the ground truth (Table 2). The only major divergence happened at a noise level of 20%, as the coefficient for $m^{2}$ had a 25% relative error. This error led to our hypothesizing that it is challenging to distinguish among the features involving $m$ (such as $m$, $m^{2}$, and $m^{3}$) due to the low range of the data. Thus, another dataset with the same properties but a larger range of values for $m$ is created. LCEN performed better on this dataset, selecting only relevant features for all noise levels tested ($\leq 30\%$) and having much lower errors in the estimated coefficients (Table 2). These experiments further highlight the robustness of LCEN and show how the range of the data can affect predictions. To clarify our design choices and the relevance of each individual part of the LCEN architecture, ablation tests are performed with this dataset (and other datasets) in Section A4 of the Appendix.
Table 2: Coefficient values and corresponding relative error to the ground truth for the “Relativistic energy” dataset at different noise levels. The first coefficient is for $m^{2}$ and should be $c^{4}=8.078$$\times$$10^{33}$ m4/s4; the second coefficient is for $m^{2}v^{2}$ and should be $c^{2}=8.988$$\times$$10^{16}$ m2/s2. The left table is for the dataset with $101\leq m<$ 10, and the right table is for the dataset with $1001\leq m<$ 100.
Noise Coefficients Error (%) 0% 8.077×$\times 10^{33}$, 8.987×$\times 10^{16}$ 0.013, 0.009 5% 8.081×$\times 10^{33}$, 8.969×$\times 10^{16}$ 0.043, 0.206 10% 8.085×$\times 10^{33}$, 8.951×$\times 10^{16}$ 0.097, 0.410 15% 8.089×$\times 10^{33}$, 8.935×$\times 10^{16}$ 0.139, 0.580 20% 6.027×$\times 10^{33}$, 8.912×$\times 10^{16}$ 25.39, 0.844    Noise Coefficients Error (%) 0% 8.078×$\times 10^{33}$, 8.987×$\times 10^{16}$ 0.001, 0.006 5% 8.078×$\times 10^{33}$, 8.986×$\times 10^{16}$ 0.005, 0.022 10% 8.078×$\times 10^{33}$, 8.984×$\times 10^{16}$ 0.009, 0.038 15% 8.079×$\times 10^{33}$, 8.983×$\times 10^{16}$ 0.013, 0.054 20% 8.079×$\times 10^{33}$, 8.981×$\times 10^{16}$ 0.017, 0.070 30% 8.080×$\times 10^{33}$, 8.978×$\times 10^{16}$ 0.025, 0.103
@@
Finally, LCEN is compared with the feature selection algorithm in ALVEN [21], which uses the same basis function expansion, but uses f-tests for feature selection. The “4th-degree, univariate polynomial” dataset is created as per Ref. [35], such that $y=X+0.5X^{2}+0.1X^{3}+0.05X^{4}+\epsilon$, 30 $X$ points are available for training, and 1,000 $X$ points are available for testing. These conditions simulate the scarcity of data potentially present in real datasets while ensuring test errors can be predicted with high confidence. Ref. [35] created four types of ALVEN models for this prediction: one that always uses degree = 4 (“unbiased model”), one that always uses degree = 2 (“biased model”), one that selects a degree between 1 and 10 based on cross-validation (“cv”), and one that selects a degree equal to 2 or 4 based on cross-validation (“cv limited order”). Ref. [35] noted that the degree 4 “unbiased model” was the best at low noise levels, but its error quickly increases, leading to the degree 2 “biased model” becoming the best for noise levels > 75 (Figure 3 of Ref. [35]; reproduced with permission here as the left subfigure of Figure 1). The model with degree equal to 2 or 4 “cv limited order” was typically very close in performance to the best model at all noise levels, whereas the model with a degree between 1 and 10 “cv” had lower performance. Ref. [35] explains these observations with the bias-variance tradeoff: at low noise levels, models should follow the ground truth as closely as possible; thus, the degree 4 “unbiased model” was the best. However, at sufficiently high noise levels, it becomes impossible to obtain enough signal to compensate for the additional degrees of freedom (variance) in a 4th degree model; thus, the degree 2 “biased model” becomes the best. The degree between 1 and 10 “cv” model had lower performance due to its greater hyperparameter variance, and the degree equal to 2 or 4 “cv limited order” model struck a balance between the “unbiased model” and the “biased model”.